# TrinityDNA: A Bio-Inspired Foundational Model for Efficient Long-Sequence DNA Modeling

### v1

@@ 4.3.1 GUE Benchmark
We evaluate our models on a comprehensive Genomic Understanding Evaluation (GUE) benchmark comprising tasks from Genomics Benchmark (zhou2023dnabert, ) and Nucleotide Transformer tasks (dalla2024nucleotide, ) by standard LoRA fine-tuning (hu2021lora, ). These include regulatory element classification, histone marker prediction, splice site annotation, and promoter/enhancer recognition tasks across various species. We refer to the best results of baselines in the original paper.
The results are shown in Table 2, comparing models DNABERT, Nucleotide Transformer (NT), DNABERT2, and our proposed TrinityDNA. Overall, TrinityDNA outperforms prior methods across many metrics (e.g., average F1 score, MCC), highlighting our model’s strong ability to capture both local motifs and long-range dependencies in genomic sequences. We observe particularly large improvements in tasks that demand recognition of extended promoter regions or higher-order structural features, aligning with our architectural design for multi-window attention and GRC-based reverse complement awareness. In summary, our results on the GUE benchmark indicate that TrinityDNA not only leverages evolutionary signals during pre-training to handle long sequences effectively but also outperforms existing models in capturing fine-scale regulatory and annotations.
Table 3: Zero-shot performance across DNA, RNA, and protein DMS tasks. All values are performance scores. Bold marks the best value per row, and underlined indicates the second best.
@@
4.3.2 Zero-shot Performance
We evaluated our two 1 B-parameter models on 19 zero-shot downstream tasks—covering DNA pathogenicity (CliVar), seven RNA DMS benchmarks and fifteen protein-fitness benchmarks across prokaryotic and eukaryotic organisms—and found that they achieve the best score on 10 of the 19 tasks, outperforming every EVO variant and the 40 M-parameter Caduceus baseline (Table 3). The prokaryote-focused TrinityMicroDNA-1B dominates the prokaryotic regime, winning 8 of 13 prokaryotic tasks and attaining the highest prokaryotic average (0.475) despite its compact size, whereas the multi-species TrinityDNA-1B excels on eukaryotic protein-fitness prediction, delivering the top Faure score and the highest eukaryotic average (0.699), even surpassing the 40 B-parameter EVO2 model (0.667). TrinityDNA also ties state-of-the-art DNA pathogenicity performance by leading on CliVar-coding and placing a close second overall. These complementary strengths underscore the benefit of evolutionary-stage-aware pre-training, and a UMAP projection of genome-level embeddings for ten representative clades (Appendix C) reveals clear taxonomic clustering, indicating that both models learn rich species-specific signals without task-specific fine-tuning.
@@ 4.3.2 Zero-shot Performance
4.3.3 CDS Annotation Benchmark
We also introduce a novel CDS Annotation Benchmark, aiming to assess long-sequence inference capabilities, practical utility for gene annotation in real-world genomes.
Data Source
From RefSeq, we collect all prokaryotic reference genomes and parse the GenBank annotation files for gene positions/types. This yields token-level labels indicating whether each token belongs to a coding sequence (CDS) with 20k sequence length and, if so, in which strand/direction it is transcribed. The detailed data statistics are described in Appendix C.1.
@@ Results
The results are shown in Table 4.3.2. While Prodigal shows strong recall performance, TrinityMicroDNA-1B delivers the best precision and F1 scores for exact matches, highlighting the model’s strong generalization capabilities across diverse datasets compared to classical pipelines.
5 Conclusion and Limitations
We present TrinityDNA, a foundational DNA model that integrates biologically inspired modules—Groove Fusion and Gated Reverse Complement—alongside a multi-scale attention mechanism to capture both local sequence signals and global genomic context effectively. Built upon an evolutionary training strategy that transitions gradually from prokaryotic to eukaryotic genomes, TrinityDNA gains strong generalization for diverse genomic prediction tasks. To further demonstrate its versatility, we introduce the CDS Annotation Benchmark, which evaluates coding sequence identification and annotation across organisms, providing a realistic standard for genome-scale annotation.
Limitations.
Although evolutionary training excels at long-sequence tasks, it can reduce performance on shorter prokaryotic sequences. Moreover, TrinityDNA has largely been validated on discriminative tasks, leaving its generative potential—such as modeling DNA–protein complexes—unexplored. Future work will refine short-sequence performance, investigate generative capabilities, and scale to larger datasets for broader genomic applications.
Appendix A Pretraining Details
A.1 Experimental Setup Details
@@ A.1.1 Hardware Configuration
A.1.2 Training Framework
@@ A.1.3 Training Optimization
Model Size Sequence Length Layers Hidden Size FFN Hidden Size Num Heads Learning Rate
Baselines 6M 8192 8 256 682 8 1.00E-03
TrinityDNA 6M 8192 8 256 682 8 1.00E-03
Baselines 40M 8192 10 576 1536 8 6.00E-04
TrinityDNA 40M 8192 10 576 1536 8 6.00E-04
Baselines 85M 8192 12 768 2048 12 5.50E-04
TrinityDNA 85M 8192 12 768 2048 12 5.50E-04
Baselines 170M 8192 24 768 2048 16 5.00E-04
TrinityDNA 170M 8192 24 768 2048 16 5.00E-04
Baselines 470M 8192 24 1280 3413 20 4.00E-04
TrinityDNA 470M 8192 24 1280 3413 20 4.00E-04
Baselines 1B 8192 24 2048 5461 32 3.00E-04
TrinityDNA 1B 8192 24 2048 5461 32 3.00E-04
TrinityDNA 1B 30720 24 2048 5461 32 2.00E-04
TrinityDNA 1B 102400 24 2048 5461 32 1.00E-04
A.2 Data
@@
We integrate raw DNA sequence data from the GTDB (Genome Taxonomy Database), IMG (Integrated Microbial Genomes), and RefSeq databases. The OpenGenome dataset was composed by sampling bacterial and archaeal genomes from the GTDB v214.1, curated prokaryotic viruses from IMG/VR v4, and plasmid sequences from IMG/PR, retaining representative genomes for each species. The Multispecies dataset comprises diverse species selected from RefSeq at the genus level, totaling 850 species and about 174 billion nucleotides, ensuring a broad coverage of evolutionary lineages. This data encompasses a wide range of genomic information, including genomic sequences and transcriptomic data. By aggregating data from multiple authoritative sources, we ensure the diversity and comprehensiveness of our dataset, enabling the model to learn a broad spectrum of DNA sequence patterns and biological features.
@@
During the pretraining phase, we adopt a masked language model (MLM) strategy. Specifically, we randomly select 15% of the tokens in the DNA sequences as masked tokens. These selected tokens are then replaced with a special $<$mask$>$ token. This process allows the model to learn to predict the original nucleotides based on the context provided by the surrounding unmasked tokens.
However, since the $<$mask$>$ token is typically not present in downstream fine-tuning tasks, we introduce a more sophisticated replacement rule to mitigate the inconsistency between pretraining and fine-tuning. Among the 15% of selected tokens: 1. With an 80% probability, a token is replaced with the $<$mask$>$ token. To further align the long-sequence training process with downstream fine-tuning, there is a 0.02 probability that the replacement ratio can range from 0 to 80%. 2. With a 10% probability, a token is replaced with a randomly chosen token from the vocabulary. 3. With a 10% probability, a token remains unchanged.
@@ A.3 Models
Appendix B Downstream Task Details
In our study, apart from the Zero-shot benchmark, all other downstream tasks are carried out in the form of LoRA (Low-rank Adaptation) fine-tuning [12]. Table 6 presents the detailed LoRA fine-tuning parameters for a variety of tasks, which is essential for understanding the specific configurations and experimental setups of each task.
B.1 LoRA Fine-tuning Parameters on GUE Benchmark [39]
@@ B.1.1 Task Descriptions
This task focuses on detecting proximal promoter regions, the critical genomic sequences that initiate transcription. Because these segments host numerous key regulatory elements, precise identification is essential for advancing our understanding of gene regulatory mechanisms and recognizing the genomic basis of various diseases. Following [7], promoter sequences are split into TATA and non-TATA based on whether they contain a TATA box motif. For each subgroup, we extract the region from 249-249- 249 to +50+50+ 50 base pairs around the transcription start site (TSS) to form the positive (promoter) class. For the negative (non-promoter) class, we randomly select equally sized sequences that (1) contain a TATA motif but lie outside promoter regions (TATA non-promoters), or (2) are formed by random substitutions (non-TATA non-promoters). We also merge both TATA and non-TATA subsets into a combined dataset labeled all.
@@
This task involves forecasting transcription factor (TF) binding sites in the human genome. TF binding sites are central to gene expression regulation; accurate prediction helps decode intricate gene regulatory networks and identify therapeutic targets. We use 690 ENCODE ChIP-seq experiments [29], covering 161 TF binding profiles in 91 cell lines. A 101-bp window around the center of each peak defines the positive (TFBS) class, while segments of the same length and matched GC content form the negative (non-TFBS) class. To avoid trivial or overly difficult tasks (e.g., F1 scores $>$ 0.95 or $<$ 0.50), we filter out such datasets and randomly pick 5 from the remaining pool.
@@
This task locates splice donor and acceptor sites in the human genome. Correct splice site identification is crucial for understanding protein diversity and the pathological effects of aberrant splicing. We use a dataset [33] of 400-bp sequences from the Ensembl GRCh38 reference genome. Following [13], we note that models often attain near-perfect results on the original set (10,000 splice donors, acceptors, and non-splice sites), which does not reflect the real difficulty of detecting non-canonical sites. To address this, we iteratively enrich the dataset with adversarial examples (previously unseen false positives) in a hold-out set, making the task substantially harder.
@@
Here, the goal is to classify species based on genomic segments. We build these datasets from viral and fungal reference genomes obtained from GenBank [2].
@@
Analogous to the human TFBS task, this variant deals with mouse genomes using Mouse ENCODE ChIP-seq data [27]. We generate negative sequences by applying dinucleotide shuffling while preserving frequency. All other settings remain consistent with the human TFBS dataset. As with the human version, we filter out tasks with extreme F1 scores and randomly sample 5 datasets from the remaining collection.
@@
This task forecasts epigenetic modifications (e.g., DNA or histone modifications) in yeast. Such marks affect gene regulation without altering the DNA sequence. We downloaded 10 datasets from [30] and split each into training, validation, and test sets with an 8:1:1 ratio.
@@
Focusing on SARS-CoV-2, this task predicts the virus’s variant type from 1000-bp genomic fragments. We gather data from EpiCoV [14] provided by GISAID, covering 9 variants: Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda, and Zeta.
@@
Appendix C Zero-shot Details
C.0.1 Datasets.
@@ Appendix C Zero-shot Details
We follow the pipeline of [21]: single-nucleotide polymorphisms (SNPs) labelled as pathogenic or benign are taken from the ClinVar release, and up to 4k nucleotide flanking windows are retrieved from the GRCh38 human reference genome to form input sequences.
@@
Seven non-coding RNA DMS benchmarks—Zhang, Pitt, Hayden, Guy, Kobori, Domingo and Andreasson—are collected exactly as in [21]. They cover diverse prokaryotic RNAs.
@@
Starting from the proteinGYM suite [23], we keep every experiment for which the wild-type codon context can be unambiguously reconstructed. This yields nine prokaryotic sets (Firnberg, Jacquier, Kelsic, Weeks, Rockah, Chen) and one human set (Sun). To broaden the eukaryotic coverage, we further add two recently released human protein scans. For cases where the original authors reported only amino-acid sequences, we rebuild nucleic-acid inputs by selecting the most frequent codon for each amino-acid according to a codon-usage table computed from all CDS regions of GRCh38.
@@
C.0.2 Zero-shot Metric Computation.
@@
Figure 8: Comparison diagrams of two examples in the CDS annotation task. The orange parts are the annotation results of the prodigal (serving as the ground truth), and the red parts are the prediction results of TrinityDNA.
C.1 CDS Benchmark
C.1.1 Data and Configuration
@@
Obtain all prokaryotic reference genomes from RefSeq, and construct token-level annotation labels based on gene localization and type information in GenBank [2] annotation files.
@@
[…]
C.1.2 Evaluation Metrics on the CDS Benchmark
@@
In line with [20], we introduce a less stringent matching criterion. Under 75% Match, a predicted CDS is counted as a TP if:
@@
Appendix D Related Work

### v2

@@ 4 Experiments
Table 2: GUE Benchmark performance comparison. Metrics are single/multi-label binary classification (MCC).
@@
Table 3: Zero-shot performance across DNA, RNA, and protein DMS tasks. ClinVar tasks are binary classification (AUC); others are RNA/protein fitness regression (Spearman).
@@
Ablation Study and Analysis
(1) Effectiveness.
Table 1 shows that GFM consistently lowers perplexity by modeling the spatial ‘groove’ features of DNA sequences. Likewise, incorporating GRC (captures reverse-complement patterns) yields a notable drop in PPL, reflecting the importance of complementary-strand information. Meanwhile, SMWA enables multi-scale context handling, trading off some computational overhead for competitive perplexity across local and longer-range dependencies. Although SMWA incurs some performance loss, it is still evident that it saves resources for long sequence training.
(2) Efficiency.
The left panel of Figure 6 contrasts token-throughput as we sweep both sequence length and micro-batch size on 1B scale. Across every setting, TrinityDNA remains clearly at the top, retaining more than 80% of its short-sequence throughput even at 64k tokens. This robustness stems from its sliding multi-window attention and optimized fused kernels, which maintain memory traffic at a nearly constant level as the context grows.
(3) Training Strategies.
The right plot in Figure 6 shows a comprehensive ablation study separating the contributions of dataset size and evolutionary training strategy. The table shows perplexity scores across different model configurations. Key finding: Models initialized with weights from prokaryotic pre-training and then fine-tuned on combined data show better performance than models trained from scratch on the combined dataset. This validates both the importance of large, diverse datasets and our EST.
Downstream Tasks
GUE Benchmark
We evaluate our models on a comprehensive Genomic Understanding Evaluation (GUE) benchmark comprising tasks from Genomics Benchmark (Zhou et al. 2023) and Nucleotide Transformer tasks (Dalla-Torre et al. 2024) by standard LoRA fine-tuning (Hu et al. 2021). For baselines, results are used in the original paper.
@@ Results
In Table 2, we compare models DNABERT, Nucleotide Transformer (NT), DNABERT2, and TrinityDNA. Overall, ours outperforms prior methods across many metrics. We observe large improvements in tasks that demand recognition of extended promoter regions or higher-order structural features, aligning with our architectural design for multi-window attention and GRC-based reverse complement awareness.
Zero-shot Performance
We evaluated our two 1B models on 19 zero-shot downstream tasks—covering DNA pathogenicity (ClinVar), seven RNA DMS benchmarks and fifteen protein-fitness benchmarks across prokaryotic and eukaryotic (Table 3). The prokaryote-focused TrinityMicroDNA-1B dominates the prokaryotic regime, winning 8 of 13 prokaryotic tasks and attaining the highest prokaryotic average (0.475), whereas the multi-species TrinityDNA-1B excels on eukaryotic protein-fitness prediction, delivering the top score and the highest eukaryotic average (0.699), even surpassing the 40B EVO2 model (0.667). TrinityDNA also excels in state-of-the-art DNA pathogenicity performance, leading in CliVar-coding and ranking second overall. These complementary strengths underscore the benefit of ETS, and a UMAP projection of genome-level embeddings for ten representative clades (Appendix D) reveals clear taxonomic clustering, indicating that both models learn rich species signals without fine-tuning.
CDS Annotation Benchmark
We also introduce a novel CDS Annotation Benchmark, aiming to assess long-sequence inference capabilities, practical utility for gene annotation in real-world genomes. From RefSeq, we collect all prokaryotic reference genomes and parse the GenBank annotation files for gene positions/types. This yields token-level labels indicating whether each token belongs to a coding sequence (CDS) with 20k sequence length and, if so, in which strand/direction it is transcribed. The detailed data statistics are described in Appendix D.
Results
The results are shown in Table 4. While Prodigal shows strong recall performance, TrinityMicroDNA-1B delivers the best precision and F1 scores for exact matches, highlighting the model’s strong generalization capabilities across diverse datasets compared to classical pipelines.
5 Conclusion
We present TrinityDNA, a foundational DNA model that integrates biologically inspired modules—Groove Fusion and Gated Reverse Complement—alongside a multi-scale attention mechanism to capture scales of genomic context effectively, built upon an ETS that transitions from prokaryotic to eukaryotic genomes. TrinityDNA gains strong generalization for diverse genomic prediction tasks. We also introduce the CDS Annotation Benchmark, which evaluates coding sequence identification across organisms and provides a realistic standard for genome-scale annotation.
Appendix A Appendix of TrinityDNA
Appendix B Experimental Setups
@@
Details for Log Influential Score Analysis
The analysis of log influential scores was designed to quantitatively measure how the influence of an input token on an output token diminishes with increasing distance between them. This helps to empirically visualize the locality bias of different model architectures.
  1. 1
    Models and Dataset: We utilized the official pretrained checkpoints for the SSM-based model Caduceus-Phage and the full-attention model DNABERT2. The analysis was conducted on the human reference genome, HG-38.
  2. 2
    Sequence Sampling: We randomly selected a contiguous sequence of 20,000 base pairs from the HG-38 genome to serve as the input for the models.
  3. 3
    Influence Calculation: The influential score is defined as the gradient of a model’s output with respect to its input, specifically $|\partial y_{t}/\partial x_{s}|$. This value quantifies the extent to which a change in the input token at position $s$ affects the model’s prediction at position $t$.
    • To compute this, we performed a forward pass with the 20k-token sequence.
    • We then used backpropagation to calculate the gradients of the output logits ($y_{t}$) with respect to the input embeddings ($x_{s}$) for a chosen target position $t$ near the end of the sequence.
    • We computed the absolute value of these gradients and took their logarithm, yielding the log influential score, $\log|\partial y_{t}/\partial x_{s}|$.
  4. 4
    Plotting: The scores were plotted against the relative distance between the input and output tokens, (ts)(t-s). The curve for each model represents the average scores over multiple runs with different randomly sampled sequences. The vertical dashed line labeled “OOM at 16k” indicates the point at which DNABERT2 failed to process the sequence due to an out-of-memory error, a known limitation of its quadratic attention mechanism.
Details for Average Attention Entropy Analysis
The experiment on average attention entropy was conducted to demonstrate the “oversmoothing” phenomenon in standard Transformer models, where attention distributions become increasingly uniform as sequence lengths grow.
  1. 1
    Model Architecture: We constructed a standard two-layer Transformer encoder with full self-attention. The model was intentionally kept simple to isolate the effect of sequence length on the attention mechanism itself, without confounding factors from more complex architectures.
  2. 2
    Datasets and Training: We trained two separate instances of this model from scratch:
    • One on a large corpus of English text from Wikipedia (Text Sequence).
    • Another on genomic data from the HG-38 reference genome (Nucleotide Sequence).
    This dual-dataset approach allowed us to compare the oversmoothing effect across different data modalities.
  3. 3
    Entropy Calculation: After training, the models were evaluated on sequences of varying lengths, from $2^{10}$ (1024) to $2^{14}$ (16384) tokens.
    • For each input sequence, we extracted the attention probability distributions, $p$, from every attention head at every position.
    • We then calculated the Shannon entropy for each distribution using the formula $H(p)=-\sum_{i}p_{i}\log_{2}p_{i}$. A higher entropy value indicates a more uniform (or ”smoother”) distribution, while a lower entropy value suggests a more focused, spiky distribution.
    • The final ”Avg. Attn Entropy” value for each sequence length shown on the y-axis is the mean entropy calculated across all attention heads and all token positions in the sequence.
  4. 4
    Plotting: The chart plots the average attention entropy as a function of sequence length. The x-axis is on a logarithmic scale ($2^{N}$) to clearly show the trend across several orders of magnitude in sequence length.
Training Framework
@@
Size Sequence Length Layers Hidden Size FFN Hidden Size Num Heads Learning Rate
6M 8192 8 256 682 8 1.00E-03
40M 8192 10 576 1536 8 6.00E-04
85M 8192 12 768 2048 12 5.50E-04
170M 8192 24 768 2048 16 5.00E-04
470M 8192 24 1280 3413 20 4.00E-04
1B 8192 24 2048 5461 32 3.00E-04
1B 30720 24 2048 5461 32 2.00E-04
1B 102400 24 2048 5461 32 1.00E-04
Data
@@
The OpenGenome dataset was composed by sampling bacterial and archaeal genomes from the GTDB v214.1, curated prokaryotic viruses from IMG/VR v4, and plasmid sequences from IMG/PR data databases.
@@
During the pretraining phase, we adopt a masked language model (MLM) strategy. Specifically, we randomly select 15% of the tokens in the DNA sequences as masked tokens. These selected tokens are then replaced with a special <<mask>> token. This process allows the model to learn to predict the original nucleotides based on the context provided by the surrounding unmasked tokens.
However, since the <<mask>> token is typically not present in downstream fine-tuning tasks, we introduce a more sophisticated replacement rule to mitigate the inconsistency between pretraining and fine-tuning. Among the 15% of selected tokens: 1. With an 80% probability, a token is replaced with the <<mask>> token. To further align the long-sequence training process with downstream fine-tuning, there is a 0.02 probability that the replacement ratio can range from 0 to 80%. 2. With a 10% probability, a token is replaced with a randomly chosen token from the vocabulary. 3. With a 10% probability, a token remains unchanged.
@@
Appendix C Downstream Task Details
In our study, apart from the Zero-shot benchmark, all other downstream tasks are carried out in the form of LoRA (Low-rank Adaptation) fine-tuning (Hu et al. 2021). Table 6 presents the detailed LoRA fine-tuning parameters for a variety of tasks, which is essential for understanding the specific configurations and experimental setups of each task.
LoRA Fine-tuning Parameters on GUE Benchmark (Zhou et al. 2023)
@@ Appendix C Downstream Task Details
This task focuses on detecting proximal promoter regions, the critical genomic sequences that initiate transcription. Because these segments host numerous key regulatory elements, precise identification is essential for advancing our understanding of gene regulatory mechanisms and recognizing the genomic basis of various diseases. Following (Dreos et al. 2013), promoter sequences are split into TATA and non-TATA based on whether they contain a TATA box motif. For each subgroup, we extract the region from 249-249 to +50+50 base pairs around the transcription start site (TSS) to form the positive (promoter) class. For the negative (non-promoter) class, we randomly select equally sized sequences that (1) contain a TATA motif but lie outside promoter regions (TATA non-promoters), or (2) are formed by random substitutions (non-TATA non-promoters). We also merge both TATA and non-TATA subsets into a combined dataset labeled all.
@@
This task involves forecasting transcription factor (TF) binding sites in the human genome. TF binding sites are central to gene expression regulation; accurate prediction helps decode intricate gene regulatory networks and identify therapeutic targets. We use 690 ENCODE ChIP-seq experiments (The ENCODE Project Consortium 2012), covering 161 TF binding profiles in 91 cell lines. A 101-bp window around the center of each peak defines the positive (TFBS) class, while segments of the same length and matched GC content form the negative (non-TFBS) class. To avoid trivial or overly difficult tasks (e.g., F1 scores >> 0.95 or << 0.50), we filter out such datasets and randomly pick 5 from the remaining pool.
@@
This task locates splice donor and acceptor sites in the human genome. Correct splice site identification is crucial for understanding protein diversity and the pathological effects of aberrant splicing. We use a dataset (Wang et al. 2019) of 400-bp sequences from the Ensembl GRCh38 reference genome. Following (Ji et al. 2021), we note that models often attain near-perfect results on the original set (10,000 splice donors, acceptors, and non-splice sites), which does not reflect the real difficulty of detecting non-canonical sites. To address this, we iteratively enrich the dataset with adversarial examples (previously unseen false positives) in a hold-out set, making the task substantially harder.
@@
Here, the goal is to classify species based on genomic segments. We build these datasets from viral and fungal reference genomes obtained from GenBank (Benson et al. 2012).
@@
Analogous to the human TFBS task, this variant deals with mouse genomes using Mouse ENCODE ChIP-seq data (Stamatoyannopoulos et al. 2012). We generate negative sequences by applying dinucleotide shuffling while preserving frequency. All other settings remain consistent with the human TFBS dataset. As with the human version, we filter out tasks with extreme F1 scores and randomly sample 5 datasets from the remaining collection.
@@
This task forecasts epigenetic modifications (e.g., DNA or histone modifications) in yeast. Such marks affect gene regulation without altering the DNA sequence. We downloaded 10 datasets from the following link: http://www.jaist.ac.jp/tran/nucleosome/members.htm and split each dataset into training, validation, and test sets at an 8:1:1 ratio.
@@
Focusing on SARS-CoV-2, this task predicts the virus’s variant type from 1000-bp genomic fragments. We gather data from EpiCoV (Khare et al. 2021) provided by GISAID, covering 9 variants: Alpha, Beta, Delta, Eta, Gamma, Iota, Kappa, Lambda, and Zeta.
@@
Appendix D Zero-shot Details
Datasets.
@@ Appendix D Zero-shot Details
We follow the pipeline of (Nguyen et al. 2024): single-nucleotide polymorphisms (SNPs) labelled as pathogenic or benign are taken from the ClinVar release, and up to 4k nucleotide flanking windows are retrieved from the GRCh38 human reference genome to form input sequences.
@@
Seven non-coding RNA DMS benchmarks—Zhang, Pitt, Hayden, Guy, Kobori, Domingo and Andreasson—are collected exactly as in (Nguyen et al. 2024). They cover diverse prokaryotic RNAs.
@@
Starting from the proteinGYM suite (Notin et al. 2023), we keep every experiment for which the wild-type codon context can be unambiguously reconstructed. This yields nine prokaryotic sets (Firnberg, Jacquier, Kelsic, Weeks, Rockah, Chen) and one human set (Sun). To broaden the eukaryotic coverage, we further add two recently released human protein scans. For cases where the original authors reported only amino-acid sequences, we rebuild nucleic-acid inputs by selecting the most frequent codon for each amino-acid according to a codon-usage table computed from all CDS regions of GRCh38.
@@
Zero-shot Metric Computation.
@@
Figure 8: Comparison diagrams of two examples in the CDS annotation task. The orange parts represent the annotation results of Prodigal (serving as the ground truth), and the red parts represent the prediction results of TrinityDNA.
CDS Benchmark
Data and Configuration
@@
Obtain all prokaryotic reference genomes from RefSeq, and construct token-level annotation labels based on gene localization and type information in GenBank (Benson et al. 2012) annotation files.
@@
We adopt Low-Rank Adaptation (LoRA) (Hu et al. 2021) to fine-tune our model on the CDS Benchmark. Specifically, we use a learning rate of $r=0.001$ and a batch size of 32. For LoRA hyperparameters, we set the low-rank dimension $\text{LoRA-r}=4$ and the scaling factor $\text{LoRA-}\alpha=32$. Furthermore, we apply LoRA to two additional MLP layers in the classification head, whose hidden dimensions are 256 and 128, respectively. By decoupling the base model parameters from the low-rank adaptation parameters, LoRA allows us to retain the expressive capacity of the underlying large model while conferring adaptability to new domains with minimal additional parameters.
Evaluation Metrics on the CDS Benchmark
@@
In line with (Nebrao et al. 2021), we introduce a less stringent matching criterion. Under 75% Match, a predicted CDS is counted as a TP if:
@@
Appendix E Related Work